OpenStudy (anonymous):

Which of the following expressions is written in scientific notation? a) 73.4 x 105 b) 0.09 × 107 c) 80 x 103 d) 4.22 x 10–3

OpenStudy (anonymous):

Scientific notation are always express in 3 sig. fig and with decimal after first figure

OpenStudy (anonymous):

no! it's d. because, there must be only one number before point(.).

OpenStudy (anonymous):

I was just looking over my answer and it wouldn't be A because of what artix said...?

OpenStudy (anonymous):

But Thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

Sorry, I have one more thing i want you to look over before i submit.. Is that okay?

OpenStudy (anonymous):

yes it's okay :)

OpenStudy (anonymous):

0.0000000708 written in scientific notation is 7.08 x 10^–8?

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

Thank You so much .. :-)

OpenStudy (anonymous):

you're welcome again :)

OpenStudy (anonymous):

Lol, Please don't kill me... I have one more question..

OpenStudy (anonymous):

Which expression represents the largest number? 40.1 x 10–6 4.1 x 10–7 0.411 x 10–7 0.04001 x 10–5

OpenStudy (anonymous):

ok. i'm here.

OpenStudy (anonymous):

answer is a.

OpenStudy (anonymous):

How is A the largest?

OpenStudy (anonymous):

because a) 40.1 x 10-6 b) 0.41 x 10-6 c) 0.0411 x 10-6 d) 0.4001 x 10-6 do you understand??

OpenStudy (anonymous):

OHH! Yes! Lol :-)

OpenStudy (anonymous):

& is 1/8000 = 8.0 x 10^3

OpenStudy (anonymous):

that's false. right one is 8.0 x 10^(-3)

OpenStudy (anonymous):

These are the answer choices .. 8.0 x 103 1.25 x 10–4 125 x 10–5 1.25 x 104

OpenStudy (anonymous):

i was wrong. i'm so sorry. answer is 1.25 x 10-4. because 1/8=0.125 and 1/8000=1/8 x 10-3 0.125 x 10-3=1.25 x 10-4 do you understand??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I'm so thankful of you.

OpenStudy (anonymous):

for

OpenStudy (anonymous):

welcome :)